Barry is 8 years older than his sister. In 3 years, he will be twice as old as she will be then. How old is each now?
Answer by stanbon(75887)

Barry is 8 years older than his sister. In 3 years, he will be twice as old as she will be then. How old is each now?
--------------
Equations::
b = s + 8
b+3 = 2(s+3)
----
Substitute for "b" and solve for "s"::
s+8 + 3 = 2s+6
s = 11-6
Sister's age = 5 yrs.
----
Solve for "b":
b = s + 8
Barry's age = 5+8 = 13 yrs.
---------------
Cheers,
Stan H.
==============